# Non-existence of Ulrich modules over Cohen-Macaulay local rings

### v1

@@ 1. Introduction
An Ulrich module over a commutative noetherian local ring $R$ is a nonzero maximum Cohen-Macaulay (MCM) $R$-module $U$ with the property that its minimal number of generators, $\nu_{R}(M)$, is equal to is Hilbert-Samuel multiplicity, $e_{R}(U)$. In [2] such modules are referred to as “maximally generated” MCM modules because the inequality $\nu_{R}(M)\leq e_{R}(M)$ holds for any MCM $R$-module $M$.
The mere existence of an Ulrich $R$-module has striking implications. For instance, it implies that Lech’s conjecture holds for $R$: For any flat local map $R\to S$ of local rings, one has $e(R)\leq e(S)$; see, for instance, [8, pg. 69], or [15, Theorem 2.11]. Moreover, if $R$ admits an Ulrich module with the additional property that its class in the Grothendieck group coincides with the class of a free module, then for any $R$-module $L$ of finite length and finite projective dimension, one has $\operatorname{length}(L)\geq e(R)$; see [12] for more general results.
@@ 4. $\bullet$
See E. Coskun’s article [3] for a nice survey on Ulrich modules and Ulrich sheaves.
Ulrich asked [20, §3] whether every Cohen-Macaulay local ring admits an Ulrich [^1] He did not refer to them as “Ulrich modules”, of course. module. Yhee [21] constructed local domains of dimension two that have no Ulrich modules. However, these examples are not Cohen-Macaulay.
@@
This result packages Theorems 4.1 and 5.1.
Our technique for producing such examples involves a relationship between Ulrich modules on $R$ and Ulrich sheaves on the projective scheme $E\coloneqq\operatorname{Proj}(\operatorname{gr}_{\mathfrak{m}}(R))$. Specifically, if $U$ is an Ulrich module, then the coherent sheaf $\mathcal{F}$ on $E$ associated to the graded module $\operatorname{gr}_{\mathfrak{m}}U$ is an Ulrich sheaf. When $R$ has dimension two, $E$ is a curve and every curve admits an Ulrich sheaf, and thus at first glance this approach seems unpromising. However, $\mathcal{F}$ extends to the blowup of $\operatorname{Spec}R$ at $\mathfrak{m}$. For the rings occurring in Theorem 1.1, we show no such “extendable” Ulrich sheaves exist on $E$. See Section 2 for details.
Our examples can be obtained by starting with an ample line bundle $\mathcal{L}$ on a projective curve $C$ such that $\Gamma(C,\mathcal{L})=0$ and the canonical pairing
@@
is a surjection for all $j\geq 2$. In this situation, the local ring $R$ obtained by localizing (or completing) the section ring $S\coloneqq\bigoplus_{j\geq 0}\Gamma(C,\mathcal{L}^{j})$ at its homogeneous maximal ideal, cannot have any Ulrich modules; see Corollary 3.2.
@@ 2. Ulrich modules and Ulrich sheaves
Let $(R,\mathfrak{m},k)$ be a noetherian local ring with maximal ideal $\mathfrak{m}$ and residue field $k$. Let $M$ be a nonzero finitely generated $R$-module. The Krull dimension and depth of $M$ are denoted $\dim_{R}M$ and $0pt_{R}M$, respectively. The $R$-module $M$ is maximal Cohen-Macaulay if it is finitely generated and $0pt_{R}M=\dim R$. We write $\nu_{R}(M)$ for the minimal number of generators of $M$, and $e_{R}(M)$ for its Hilbert-Samuel multiplicity with respect to the ideal $\mathfrak{m}$.
@@
Any maximal Cohen-Macaulay $R$-module $M$ satisfies $\nu_{R}(M)\leq e_{R}(M)$; see [20, Section 3]. When $e_{R}(M)=\nu_{R}(M)$ and $M$ is nonzero it is said to be an Ulrich module.
@@
    $\operatorname{H}^{i}(X,\mathcal{F}(j))=0$ except when $i=0$, $j\geq 0$ or $i=\dim(X)$, $j\leq-\dim(X)-1$.
@@
If $\dim(X)=1$, then $\mathcal{F}$ is an Ulrich sheaf if and only if $\mathcal{F}(-1)$ has no cohomology and $\mathcal{F}\neq 0$. Thus the existence of an Ulrich sheaf for $(X,\mathcal{O}_{X}(1))$ is equivalent to the existence of a nonzero coherent sheaf on $X$ with no cohomology, and the latter property does not depend on the choice of $\mathcal{O}_{X}(1)$.
@@
Then $\mathscr{R}_{\mathfrak{m}}(R)$ is a graded $R$-algebra, called the Rees ring of $R$ with respect $\mathfrak{m}$, and $\mathscr{R}_{\mathfrak{m}}(M)$ is a finitely generated graded $\mathscr{R}_{\mathfrak{m}}(R)$-module. The canonical map $R\to\mathscr{R}_{\mathfrak{m}}(R)$ induces the morphism
@@
Thus $\operatorname{gr}_{\mathfrak{m}}(R)$ is a standard graded $k$-algebra and the $\operatorname{gr}_{\mathfrak{m}}(R)$-module $\operatorname{gr}_{\mathfrak{m}}(M)$ is finitely generated over $\operatorname{gr}_{\mathfrak{m}}(R)$ by ${\operatorname{gr}_{\mathfrak{m}}(M)}_{0}$.
@@
[…]
@@
The multiplicity and minimal number of generators of $\operatorname{gr}_{\mathfrak{m}}(U)$ as a module over $\operatorname{gr}_{\mathfrak{m}}(R)$ coincide with the corresponding invariants of the $R$-module $U$. Thus the crucial point is that $\operatorname{gr}_{\mathfrak{m}}(U)$ is maximal Cohen-Macaulay. We can assume the residue field of $R$ is infinite, and hence that there exists an $U$-regular sequence, say $\boldsymbol{x}\coloneqq x_{1},\dots,x_{d}$, such that $\boldsymbol{x}U=\mathfrak{m}U$. Since $\boldsymbol{x}$ is $U$-regular, one has an isomorphism of $\operatorname{gr}_{(\boldsymbol{x})}(R)$-modules
@@
Here $\overline{x}_{i}$ denotes the residue class of $x_{i}$ in $(\boldsymbol{x})/(\boldsymbol{x})^{2}$. Therefore $\operatorname{gr}_{(\boldsymbol{x})}(U)$ is a maximal Cohen-Macaulay module over $\operatorname{gr}_{(\boldsymbol{x})}(R)$. It remains to observe that this module is isomorphic to $\operatorname{gr}_{\mathfrak{m}}(U)$, since $\boldsymbol{x}U=\mathfrak{m}U$.
It follows that $j^{*}\mathcal{F}$, which is the coherent sheaf associated to $\operatorname{gr}_{\mathfrak{m}}(U)$, is an Ulrich sheaf; see [4, Proposition 2.1]. The coherent sheaf $\mathcal{F}(1)$ is the sheaf associated to $\mathscr{R}_{\mathfrak{m}}(\mathfrak{m}U)$ and the map $\mathcal{F}(1)\to\mathcal{F}$ induced by $t\colon\mathcal{O}_{B}(1)\to\mathcal{O}_{B}$, where $B=\operatorname{Bl}_{\mathfrak{m}}(R)$, arises from the canonical inclusion
@@ 3. Criteria for non-existence
Assume to the contrary that there exists an Ulrich $R$-module $U$. Let $\mathcal{F}$ be the strict transform of $U$ and $\mathcal{U}$ the sheaf on $E$ associated to $\operatorname{gr}_{\mathfrak{m}}(U)$. We may identify $\mathcal{U}$ with $\mathcal{F}/{\mathcal{I}}^{n}\mathcal{F}$. Setting ${\mathcal{U}}_{j}$ to be the coherent sheaf on $E$ given by ${\mathcal{I}}^{j}\mathcal{F}/{\mathcal{I}}^{n}\mathcal{F}$, we obtain a filtration
@@
is strictly increasing in $j$. Indeed $\chi(\mathcal{G}\otimes\mathcal{L}^{j})$ is a polynomial function in $j$ of degree at most one, by  [19, Lemma 33.4.5.1], and since $\mathcal{L}$ is ample and $\dim\mathcal{G}=1$, this polynomial must be linear with positive slope. See [6, Propositions 23.78, 23.79] for more precise results.
@@
Since $\mathcal{U}$ is an Ulrich sheaf, ${\mathcal{U}}(-1)$ has no cohomology; see Remark 2.1. Combined with the fact that $\dim W=1$ one gets
@@
[…]
@@
Here is one source of rings satisfying the hypotheses of the preceding result.
@@
Let $V$ be a projective $k$-scheme with $\dim V\geq 1$, let $\mathcal{L}$ be an ample line bundle on $V$, and $S=\operatorname{S}(V,\mathcal{L})$ the section ring of $\mathcal{L}$. Suppose there is an integer $a\geq 2$ so that the following conditions hold:
@@
The second one holds as $\oplus_{j\geqslant 0}S_{\geqslant ja}t^{j}$ is the $a$-th Veronese subring of $\oplus_{j\geqslant 0}S_{\geqslant j}t^{j}$, and the last isomorphism holds because $\mathcal{L}$ is ample; see [7, Section 8.7.3] or [11, Paragraph 6.2.1]. These isomorphisms are compatible with projections to $\operatorname{Spec}S$.
@@
The exceptional fiber $E$ of the blowup $\operatorname{Bl}_{\mathfrak{m}}(S)\to\operatorname{Spec}S$ is the subscheme of $\operatorname{Bl}_{\mathfrak{m}}(S)$ cut out by $\mathcal{L}^{a}$, that is to say that
$E\cong\operatorname{\mathbf{Spec}}\left(\frac{\mathrm{Sym}(\mathcal{L})}{\mathcal{L}^{a}\,\mathrm{Sym}(\mathcal{L})}\right)\,.$
@@
Conditions (1) and (2) imply that $S$ is generated as a $k$-algebra by its components in degrees $[a,2a-1]$. However the latter condition is weaker: the $k$-algebra $k[x,y]$ with $|x|=2$ and $|y|=3$ is generated by its components in degrees $[2,3]$, but the map $S_{2}\otimes_{k}S_{4}\to S_{6}$ is not surjective for $y^{2}$ is not in the image.
Corollary 3.2 brings up the question: When is a noetherian graded ring over a field the section ring of an ample line bundle on a projective scheme? Here is one answer; it is extrapolates from [11, Proposition 2.1], where the result is stated for normal domains. The argument does not require these conditions, and we sketch a proof, to convince ourselves.
@@
Let $k$ be a field and $S$ a noetherian $\mathbb{N}$-graded $k$-algebra with $S_{0}=k$ and $0ptS\geq 2$. Then $S$ is the section ring of an ample line bundle on a projective scheme over $k$ if and only if there exist homogeneous elements $x_{0},\dots,x_{d}$ in $S$ such that $\sqrt{(x_{0},\dots,x_{d})}=S_{\geqslant 1}$ and the $\mathbb{Z}$-graded rings $S_{x_{i}}$ has a unit of degree $1$.
@@
Suppose $S=\operatorname{S}(X,\mathcal{L})$. Choose an integer $m\gg 0$ such that $\mathcal{L}^{m}$ and $\mathcal{L}^{m+1}$ are very ample (see [9, Theorem II, 7.6] and [14, Theorem 1.2.6]) and let $s_{1},\dots,s_{r}$ and $t_{1},\dots,t_{l}$ be bases of $\Gamma(X,\mathcal{L}^{m})$ and $\Gamma(X,\mathcal{L}^{m+1})$, respectively. One can check that the radical of the ideal generated by $x_{ij}=s_{i}t_{j}$, for $1\leq i\leq r$ and $1\leq j\leq l$, is $S_{\geqslant 1}$ and that $t_{j}^{2}/x_{ij}$ in $S_{x_{ij}}$ is a unit of degree 1.
For the converse, by computing global sections using Čech cover associated to the sequence $\boldsymbol{x}=(x_{0},\dots,x_{d})$, one gets an exact sequence
@@
$\operatorname{Proj}S=D(x_{0})\cup\cdots\cup D(x_{d})$
such that one has an isomorphism of $\mathbb{Z}$-graded rings $S_{x_{i}}\cong[S_{x_{i}}]_{0}[t^{\pm 1}]$, where $|t|=1$. Using this it is easy to check that the natural map
@@
Here is another special case of Theorem 3.1.
Corollary 3.5.
Proof.
Our interest in the formulation above is that it applies also when $R$ is not itself a section ring. To illustrate this, we revisit the example constructed by Yhee [21].
Yhee’s example revisited
Let $k$ be a field, $S=k[\boldsymbol{x}]$ the polynomial ring over $k$ in indeterminates $\boldsymbol{x}=x_{1},\dots,x_{d}$, and set $\mathfrak{m}=(\boldsymbol{x})$. Fix an integer $n\geq 2$, set $J=\mathfrak{m}^{n}$ and $R=k\oplus J$, viewed as a subring of $S$. Then $R$ cannot be a section ring, for it is not $S_{2}$.
@@
The second isomorphism holds as $\mathscr{R}_{J}(S)$ is the $n$th Veronese subring of $\mathscr{R}_{\mathfrak{m}}(S)$; this holds because $J=\mathfrak{m}^{n}$. In particular $\operatorname{Bl}_{J}(R)$ is regular. The exceptional fiber of the blowup is $E=\operatorname{Proj}(\operatorname{gr}_{J}(R))$ where

### v2

@@ 1. Introduction
This work concerns the question of existence of certain objects over a commutative noetherian local ring $R$. A prime example of such an $R$ is the local ring of functions at a point on an algebraic variety. In this case the structural properties of $R$ reflect the geometric properties of the variety at that point. For simplicity of exposition, we assume the ring $R$ is a domain, which in the geometric context corresponds to varieties that are irreducible at the point represented by the variety. As for groups, one can study $R$ through its representations, which in this context are called modules. In what follows we focus on a class of special class of modules called maximal Cohen-Macaulay modules, and assume that $R$ itself is Cohen-Macaulay; this ensures that there are many maximal Cohen-Macaulay modules. When $R$ is a regular local ring (which is the analogue of a smooth point on a variety), maximal Cohen-Macaulay modules are all free, and in fact this property characterizes regularity. So the failure of maximal Cohen-Macaulay modules to be free is one measure of the singularity of the ring $R$.
Two of the most basic invariants attached to any module $M$ are the minimal number of elements needed to generate it, which is denoted $\nu_{R}(M)$, and the largest cardinality of a linearly independent set, which is called its rank and denoted $\mathrm{rank}_{R}(M)$. Clearly $\mathrm{rank}_{R}(M)\leq\nu_{R}(M)$, but in general there is no a priori bound on the gap between these invariants. However, when $M$ is a maximal Cohen-Macaulay module, Ulrich [22] proved that there is an inequality $\nu_{R}(M)\leq e(R)\cdot\mathrm{rank}_{R}(M)$. Here $e(R)$ is the Hilbert-Samuel multiplicity of $R$, which is one measure of singularity of $R$. For instance, regular local rings $R$ are characterized by the property that $e(R)=1$.
Ulrich asked whether there always exist a module $M$ for which $\nu_{R}(M)=e(R)\cdot\mathrm{rank}_{R}(M)$; such modules were subsequently named Ulrich modules. It follows from the discussion above that $R$ itself is an Ulrich module precisely when $R$ is regular. Inspired by Ulrich’s work, in 2003 Eisenbud and Schreyer [4] introduced an analogous notion for projective varieties in algebraic geometry, which they called Ulrich sheaves. When $R$ has an Ulrich module, one can often reduce problems over $R$ to the case of regular rings, where they are more tractable. In the almost four decades since Ulrich modules were introduced, it was discovered that many open conjectures in commutative algebra and algebra geometry are simple consequences of the existence of Ulrich modules and sheaves. In the meantime, Ulrich modules and sheaves have been constructed in many special cases, under various algebraic and geometric assumptions.
In this paper we construct Cohen-Macaulay rings of dimension two and higher, some of which are even Gorenstein and normal, and others that are complete intersections, that possess no Ulrich modules. Our technique for producing such examples has a geometric flavor, as it involves a relationship between Ulrich modules on $R$ and Ulrich sheaves on a projective variety naturally associated to $R$, namely, the exceptional fiber of the blow-up of $R$ at its maximal ideal. Ulrich sheaves that arise thus are special in that they can be extended to the entire blow-up.
While this paper settles Ulrich’s question in the negative on the algebraic side, the existence of Ulrich sheaves remains an open and interesting question.
A more detailed introduction
An Ulrich module over an arbitrary commutative noetherian local ring $R$ is a nonzero maximal Cohen-Macaulay $R$-module $U$ with the property that its minimal number of generators, $\nu_{R}(U)$, is equal to its Hilbert-Samuel multiplicity, $e_{R}(U)$. In [2] such modules are referred to as “maximally generated” maximal Cohen-Macaulay modules because the inequality $\nu_{R}(M)\leq e_{R}(M)$ holds for any maximal Cohen-Macaulay $R$-module $M$.
The mere existence of an Ulrich $R$-module has striking implications. For instance, it implies that Lech’s conjecture holds for $R$: For any flat local map $R\to S$ of local rings, one has $e(R)\leq e(S)$; see [9, pg. 69], or [17, Theorem 2.11]. Moreover, if $R$ admits an Ulrich module with the additional property that its class in the Grothendieck group coincides with the class of a free module, then for any $R$-module $L$ of finite length and finite projective dimension, one has $\operatorname{length}_{R}(L)\geq e(R)$; see [14] for more general results.
@@ 4. $\bullet$
See Emre Coskun’s article [3] for a nice survey on Ulrich modules and Ulrich sheaves.
Ulrich asked [22, §3] whether every Cohen-Macaulay local ring admits an Ulrich [^1] He did not refer to them as “Ulrich modules”. module. Yhee [23] constructed local domains of dimension two that have no Ulrich modules. However, these examples are not Cohen-Macaulay.
@@
This result packages Theorems 4.1 and 5.1. We do not know if the rings we construct admit lim Ulrich sequences, in the sense of [17].
Let $B$ be the blowup of $\operatorname{Spec}(R)$ at its closed point and $E=\operatorname{Proj}(\operatorname{gr}_{\mathfrak{m}}(R))$ the exceptional fiber. As mentioned, the existence of an Ulrich module on $R$ leads to an Ulrich sheaf on $E$ that extends to $B$. Our technique is to find examples where there are no extendable Ulrich sheaves on $E$. All of our examples can be obtained by starting with an ample line bundle $\mathcal{L}$ on a projective curve $C$ such that $\Gamma(C,\mathcal{L})=0$ and the canonical pairing
@@
is a surjection for all $j\geq 2$. In this situation, the local ring $R$ obtained by localizing (or completing) the section ring $\bigoplus_{j\geq 0}\Gamma(C,\mathcal{L}^{j})$ at its homogeneous maximal ideal, cannot have any Ulrich modules; see Corollary 3.3.
@@ 2. Ulrich modules and Ulrich sheaves
Let $(R,\mathfrak{m},k)$ be a noetherian local ring with maximal ideal $\mathfrak{m}$ and residue field $k$. Let $M$ be a nonzero finitely generated $R$-module. The Krull dimension and depth of $M$ are denoted $\dim_{R}M$ and $0pt_{R}(M)$, respectively. The $R$-module $M$ is maximal Cohen-Macaulay if it is finitely generated and $0pt_{R}(M)=\dim R$. We write $\nu_{R}(M)$ for the minimal number of generators of $M$, and $e_{R}(M)$ for its Hilbert-Samuel multiplicity with respect to the ideal $\mathfrak{m}$.
@@
Any maximal Cohen-Macaulay $R$-module $M$ satisfies $\nu_{R}(M)\leq e_{R}(M)$; see [22, Section 3]. When $e_{R}(M)=\nu_{R}(M)$ and $M$ is nonzero it is said to be an Ulrich module.
@@
    $\operatorname{H}^{i}(X,\mathcal{F}(j))=0$ except when $i=0$, $j\geq 0$ or $i=\dim X$, $j\leq-\dim X-1$.
@@
If $\dim X=1$, then $\mathcal{F}$ is an Ulrich sheaf if and only if $\operatorname{H}^{i}(X,\mathcal{F}(-1))=0$ for all $i$ and $\mathcal{F}\neq 0$. Thus the existence of an Ulrich sheaf for $(X,\mathcal{O}_{X}(1))$ is equivalent to the existence of a nonzero coherent sheaf on $X$ with no cohomology, and the latter property does not depend on the choice of $\mathcal{O}_{X}(1)$.
@@
Then $\mathscr{R}_{\mathfrak{m}}(R)$ is a finitely generated graded $R$-algebra, called the Rees ring of $R$ with respect $\mathfrak{m}$, and $\mathscr{R}_{\mathfrak{m}}(M)$ is a finitely generated graded $\mathscr{R}_{\mathfrak{m}}(R)$-module. The canonical map $R\to\mathscr{R}_{\mathfrak{m}}(R)$ induces the morphism
@@
Thus $\operatorname{gr}_{\mathfrak{m}}(R)$ is a finitely generated standard graded $k$-algebra and the $\operatorname{gr}_{\mathfrak{m}}(R)$-module $\operatorname{gr}_{\mathfrak{m}}(M)$ is finitely generated by ${\operatorname{gr}_{\mathfrak{m}}(M)}_{0}$.
@@
${E}$${\operatorname{Bl}_{\mathfrak{m}}(R)}$${\operatorname{Spec}k}$${\operatorname{Spec}R}$$\scriptstyle{q}$$\scriptstyle{j}$$\scriptstyle{p}$
@@
The multiplicity and minimal number of generators of $\operatorname{gr}_{\mathfrak{m}}(U)$ as a module over $\operatorname{gr}_{\mathfrak{m}}(R)$ coincide with the corresponding invariants of the $R$-module $U$. Thus the crucial point is that $\operatorname{gr}_{\mathfrak{m}}(U)$ is maximal Cohen-Macaulay. We can assume the residue field of $R$ is infinite, and hence that there exists a $U$-regular sequence, say $\boldsymbol{x}\coloneqq x_{1},\dots,x_{d}$, such that $\boldsymbol{x}U=\mathfrak{m}U$. Since $\boldsymbol{x}$ is $U$-regular, one has an isomorphism of $\operatorname{gr}_{(\boldsymbol{x})}(R)$-modules
@@
Here $\overline{x}_{i}$ denotes the residue class of $x_{i}$ in $(\boldsymbol{x})/(\boldsymbol{x})^{2}$. Therefore $\operatorname{gr}_{(\boldsymbol{x})}(U)$ is a maximal Cohen-Macaulay module over $\operatorname{gr}_{(\boldsymbol{x})}(R)$. It remains to observe that $\operatorname{gr}_{(\boldsymbol{x})}(U)$ is isomorphic to $\operatorname{gr}_{\mathfrak{m}}(U)$, since $\boldsymbol{x}U=\mathfrak{m}U$.
It follows that $j^{*}\mathcal{F}$, which is the coherent sheaf associated to $\operatorname{gr}_{\mathfrak{m}}(U)$, is an Ulrich sheaf; see [4, Proposition 2.1]. The coherent sheaf $\mathcal{F}(1)$ is the sheaf associated to $\mathscr{R}_{\mathfrak{m}}(\mathfrak{m}U)$ and the map $\mathcal{F}(1)\to\mathcal{F}$ induced by the natural map $\mathcal{O}_{B}(1)\to\mathcal{O}_{B}$, where $B=\operatorname{Bl}_{\mathfrak{m}}(R)$, arises from the canonical inclusion
@@ 3. Criteria for non-existence
Assume to the contrary that there exists an Ulrich $R$-module $U$. Let $\mathcal{F}$ be the strict transform of $U$ and $\mathcal{U}$ the sheaf on $E$ associated to $\operatorname{gr}_{\mathfrak{m}}(U)$. We may identify $\mathcal{U}$ with $\mathcal{F}/{\mathcal{I}}^{n}\mathcal{F}$. Setting ${\mathcal{U}}_{i}$ to be the coherent sheaf on $E$ given by ${\mathcal{I}}^{i}\mathcal{F}/{\mathcal{I}}^{n}\mathcal{F}$, we obtain a filtration
@@
is strictly increasing in $j$. Indeed $\chi(\mathcal{G}\otimes\mathcal{L}^{j})$ is a polynomial function in $j$ of degree at most one, by  [21, Lemma 33.4.5.1], and since $\mathcal{L}$ is ample and $\dim\mathcal{G}=1$, this polynomial must be linear with positive slope. See [6, Propositions 23.78, 23.79] for more precise results.
@@
Since $\mathcal{U}$ is an Ulrich sheaf, ${\mathcal{U}}(-1)$ has no cohomology. Combined with the fact that $\dim W=1$ one gets
@@
${\mathscr{R}_{\mathfrak{m}}(R)}$${\mathscr{R}_{\mathfrak{m}R^{\prime}}(R^{\prime})}$${\operatorname{gr}_{\mathfrak{m}}(R)}$${\operatorname{gr}_{\mathfrak{m}R^{\prime}}(R^{\prime})}$
@@
Here is a special case of Theorem 3.1.
Corollary 3.2.
Proof.
Our interest in the formulation above is that it applies also when $R$ is not a section ring; see Remark 3.6 where we revisit the example constructed by Yhee [23].
Next we identify another source of rings satisfying the hypotheses of Theorem 3.1.
@@
Let $V$ be a projective $k$-scheme with $\dim V\geq 1$, let $\mathcal{L}$ be an ample line bundle on $V$, and set $S=\operatorname{S}(V,\mathcal{L})$. Suppose there is an integer $a\geq 2$ so that the following conditions hold:
@@
The second one holds as $\oplus_{j\geqslant 0}S_{\geqslant ja}t^{j}$ is the $a$-th Veronese subring of $\oplus_{j\geqslant 0}S_{\geqslant j}t^{j}$, and the last isomorphism holds because $\mathcal{L}$ is ample; see [8, Section 8.7.3] or [12, Paragraph 6.2.1]. These isomorphisms are compatible with projections to $\operatorname{Spec}S$.
@@
The exceptional fiber $E$ of the blowup $\operatorname{Bl}_{\mathfrak{m}}(S)\to\operatorname{Spec}S$ is the subscheme of $\operatorname{Bl}_{\mathfrak{m}}(S)$ cut out by $\mathcal{L}^{a}$, that is to say
$E\cong\operatorname{\mathbf{Spec}}_{V}\left(\frac{\mathrm{Sym}(\mathcal{L})}{\mathcal{L}^{a}\,\mathrm{Sym}(\mathcal{L})}\right)\,.$
@@
Conditions (1) and (2) imply that $S$ is generated as a $k$-algebra by its components in degrees $[a,2a-1]$. However the latter condition is weaker: the $k$-algebra $k[x,y]$ with $|x|=2$ and $|y|=3$ is generated by its components in degrees $[2,3]$, but the map $S_{2}\otimes_{k}S_{4}\to S_{6}$ is not surjective because $y^{2}$ is not in the image.
Corollary 3.3 brings up the question: When is a noetherian graded ring over a field the section ring of an ample line bundle on a projective scheme? Here is one answer; it is extrapolates from [13, Proposition 2.1], where the result is stated for normal domains. The argument does not require these conditions, and we sketch a proof, to convince ourselves.
@@
Let $k$ be a field and $S$ a noetherian $\mathbb{N}$-graded $k$-algebra with $S_{0}=k$ and $0ptS\geq 2$. Then $S$ is the section ring of an ample line bundle on a projective scheme over $k$ if and only if there exist homogeneous elements $x_{0},\dots,x_{n}$ in $S$ such that $\sqrt{(x_{0},\dots,x_{n})}=S_{\geqslant 1}$ and the $\mathbb{Z}$-graded rings $S_{x_{i}}$ have a unit of degree $1$.
@@
Suppose $S=\operatorname{S}(X,\mathcal{L})$. Choose an integer $m\gg 0$ such that $\mathcal{L}^{m}$ and $\mathcal{L}^{m+1}$ are very ample (see [10, Theorem II, 7.6] and [16, Theorem 1.2.6]) and let $s_{1},\dots,s_{r}$ and $t_{1},\dots,t_{l}$ be bases of $\Gamma(X,\mathcal{L}^{m})$ and $\Gamma(X,\mathcal{L}^{m+1})$, respectively. One can check that the radical of the ideal generated by $x_{ij}=s_{i}t_{j}$, for $1\leq i\leq r$ and $1\leq j\leq l$, is $S_{\geqslant 1}$ and that $t_{j}^{2}/x_{ij}$ in $S_{x_{ij}}$ is a unit of degree 1.
For the converse, by computing global sections using the Čech cover associated to the sequence $\boldsymbol{x}=x_{0},\dots,x_{n}$, one gets an exact sequence
@@
$\operatorname{Proj}S=D(x_{0})\cup\cdots\cup D(x_{n})$
such that one has an isomorphism of $\mathbb{Z}$-graded rings $S_{x_{i}}\cong[S_{x_{i}}]_{0}[t,t^{-1}]$, where $t$ has degree $1$. Using this it is easy to check that the natural map
@@
Remark 3.6.
(Yhee’s example revisited) Let $k$ be a field, $S=k[\boldsymbol{x}]$ the polynomial ring over $k$ in indeterminates $\boldsymbol{x}=x_{1},\dots,x_{d}$, and set $\mathfrak{m}=(\boldsymbol{x})$. Fix an integer $n\geq 2$, set $J=\mathfrak{m}^{n}$ and $R=k\oplus J$, viewed as a subring of $S$. Then $R$ is not a section ring.
@@
The second isomorphism holds as $\mathscr{R}_{J}(S)$ is the $n$th Veronese subring of $\mathscr{R}_{\mathfrak{m}}(S)$, since $J=\mathfrak{m}^{n}$. In particular $\operatorname{Bl}_{J}(R)$ is regular. The exceptional fiber of the blowup is $E=\operatorname{Proj}(\operatorname{gr}_{J}(R))$ where